# Extreme Precipitation Nowcasting using Multi-Task Latent Diffusion Models

### v2

@@ IV-C Diffusion Model Training and Sampling Overview
The diffusion model training involves several steps. Firstly, initial data is sampled from the real data distribution. Then, a time step in the diffusion process is randomly selected, and noise $\epsilon$ is sampled from the standard normal distribution. The difference between the model prediction noise and the actual noise is calculated and minimized. Model parameters are updated via gradient descent, and this process is repeated until the model converges, enhancing the model’s ability to recover the original data from noise. The training is carried out using the PyTorch 1.3.1 framework on 10 A6000 GPU devices under Ubuntu 22.04. In the inference phase, the DDPM sampling method is used on a single A800 GPU. It takes approximately 480 seconds to sample a batch of data, which means generating 16 frames of prediction results. For the diffusion model and denoising network, SSLDM and MTLDM have the same structure. The latent space encoder $D_{0}$ and the conditional encoder are also identical. The distinction between the two lies in the need to train different decoders for different decomposed sub-images.
@@ IV-D Evaluation based on Metrics
TABLE II: Comparison of precipitation CSI (↑) on the MRMS dataset. ↑ means the larger the better
No PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) MTLDM PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) MTLDM
@@
1 0.5292 0.5798 0.4686 0.4692 0.6243 0.6262 0.2194 0.2621 0.1489 0.1355 0.2796 0.2728
2 0.4912 0.5144 0.4448 0.4379 0.5670 0.5728 0.1858 0.1949 0.1135 0.1094 0.2066 0.2157
3 0.4455 0.4804 0.4249 0.4153 0.5272 0.5359 0.1499 0.1585 0.0905 0.0887 0.1654 0.1758
4 0.4146 0.4407 0.4083 0.3963 0.4943 0.5016 0.1326 0.1292 0.0767 0.0792 0.1422 0.1552
5 0.3875 0.4173 0.3931 0.3798 0.4677 0.4752 0.1132 0.1090 0.0644 0.0672 0.1261 0.1366
6 0.3644 0.3958 0.3805 0.3629 0.4439 0.4536 0.1027 0.0947 0.0576 0.0608 0.1125 0.1224
7 0.3420 0.3739 0.3648 0.3461 0.4229 0.4311 0.0916 0.0827 0.0484 0.0520 0.0990 0.1117
8 0.3234 0.3530 0.3524 0.3310 0.4015 0.4098 0.0826 0.0727 0.0437 0.0471 0.0866 0.0978
9 0.3065 0.3358 0.3408 0.3191 0.3831 0.3916 0.0748 0.0642 0.0396 0.0424 0.0763 0.0888
10 0.2930 0.3210 0.3306 0.3081 0.3690 0.3766 0.0697 0.0573 0.0351 0.0382 0.0695 0.080
11 0.2794 0.3074 0.3199 0.2959 0.3555 0.3628 0.0643 0.0519 0.0319 0.0345 0.0642 0.0753
12 0.2668 0.2945 0.3101 0.2865 0.3418 0.3494 0.0584 0.0469 0.0298 0.0327 0.0586 0.0682
13 0.2570 0.2830 0.3015 0.2784 0.3318 0.3388 0.0542 0.0437 0.0276 0.0306 0.0562 0.0650
14 0.2467 0.2719 0.2933 0.2692 0.3198 0.3271 0.0522 0.0404 0.0256 0.0273 0.0524 0.0601
15 0.2377 0.2617 0.2858 0.2618 0.3092 0.3176 0.0494 0.0369 0.0238 0.0260 0.0493 0.0564
16 0.2289 0.2512 0.2773 0.2515 0.2999 0.3072 0.0462 0.0330 0.0242 0.0260 0.0469 0.0537
No PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) MTLDM PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) MTLDM
@@
1 0.1880 0.2271 0.1271 0.1163 0.2442 0.2355 0.0810 0.1035 0.0411 0.0310 0.1048 0.0979
2 0.1565 0.1647 0.0977 0.0950 0.1744 0.1818 0.0636 0.0641 0.0318 0.0168 0.0660 0.0711
3 0.1257 0.1329 0.0787 0.0771 0.1386 0.1471 0.0430 0.0428 0.0234 0.0092 0.0457 0.0530
4 0.1103 0.1077 0.0677 0.0700 0.1187 0.1273 0.0338 0.0311 0.0184 0.0073 0.0366 0.0434
5 0.0946 0.0916 0.0568 0.0597 0.1051 0.1138 0.0274 0.0238 0.0144 0.0054 0.0286 0.0365
6 0.0854 0.0801 0.0511 0.0539 0.0942 0.1030 0.0250 0.0171 0.0133 0.0049 0.0243 0.0300
7 0.0760 0.0699 0.0438 0.0472 0.0828 0.0919 0.0208 0.0135 0.0107 0.0028 0.0189 0.0246
8 0.0690 0.0615 0.0399 0.0429 0.0731 0.0819 0.0189 0.0119 0.0076 0.0018 0.0152 0.0203
9 0.0628 0.0546 0.0360 0.0384 0.0642 0.0747 0.0151 0.0106 0.0073 0.0025 0.0122 0.0175
10 0.0586 0.0486 0.0320 0.0350 0.0599 0.0681 0.0131 0.0080 0.0063 0.0019 0.0091 0.0143
11 0.0540 0.0444 0.0290 0.0313 0.0549 0.0641 0.0122 0.0067 0.0060 0.0014 0.0070 0.0117
12 0.0497 0.0404 0.0276 0.0302 0.0506 0.0591 0.0109 0.0064 0.0048 0.0015 0.0056 0.0105
13 0.0462 0.0379 0.0259 0.0285 0.0493 0.0560 0.0095 0.0052 0.0038 0.0011 0.0059 0.0086
14 0.0448 0.0348 0.0238 0.0257 0.0462 0.0520 0.0093 0.0050 0.0038 0.0008 0.0048 0.0079
15 0.0422 0.0318 0.0221 0.0241 0.0432 0.0487 0.0079 0.0051 0.0039 0.0009 0.0055 0.0076
16 0.0393 0.0282 0.0222 0.0239 0.0407 0.0467 0.0071 0.0048 0.0041 0.0013 0.0048 0.0078
TABLE IV: Comparison of precipitation CSI (↑) on the SW dataset. ↑ means the larger the better
No PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) MTLDM PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) MTLDM
@@
1 0.4235 0.3773 0.2889 0.2655 0.3934 0.4663 0.1882 0.1657 0.0928 0.0740 0.1602 0.2642
2 0.3765 0.2878 0.2550 0.2443 0.3497 0.4148 0.1557 0.1132 0.0729 0.0634 0.1331 0.2067
3 0.3386 0.2724 0.2322 0.2267 0.3155 0.3737 0.1320 0.1105 0.0602 0.0559 0.1134 0.1649
4 0.3084 0.2368 0.2137 0.2117 0.2855 0.3318 0.1135 0.0871 0.0512 0.0507 0.0965 0.1337
5 0.2821 0.2090 0.1962 0.1971 0.2588 0.3029 0.0993 0.0647 0.0438 0.0423 0.0829 0.1114
6 0.2596 0.1859 0.1821 0.1840 0.2365 0.2761 0.0865 0.0527 0.0375 0.0379 0.0717 0.0954
7 0.2420 0.1791 0.1701 0.1711 0.2188 0.2564 0.0787 0.0528 0.0330 0.0331 0.0655 0.0836
8 0.2265 0.1634 0.1602 0.1614 0.2041 0.2371 0.0710 0.0441 0.0291 0.0296 0.0576 0.0728
9 0.2116 0.1586 0.1508 0.1518 0.1898 0.2206 0.0639 0.0393 0.0261 0.0259 0.0518 0.0639
10 0.1991 0.1507 0.1414 0.1437 0.1781 0.2055 0.0593 0.0379 0.0240 0.0226 0.0481 0.0561
11 0.1879 0.1408 0.1342 0.1356 0.1672 0.1934 0.0544 0.0362 0.0210 0.0218 0.0433 0.0501
12 0.1778 0.1279 0.1260 0.1286 0.1573 0.1816 0.0499 0.0359 0.0183 0.0194 0.0403 0.0446
13 0.1685 0.1371 0.1191 0.1210 0.1489 0.1712 0.0458 0.0382 0.0164 0.0170 0.0383 0.0404
14 0.1594 0.1142 0.1117 0.1136 0.1418 0.1619 0.0409 0.0320 0.0151 0.0152 0.0346 0.0367
15 0.1519 0.1164 0.1059 0.1069 0.1350 0.1546 0.0376 0.0258 0.0149 0.0141 0.0323 0.0339
16 0.1444 0.1227 0.1001 0.1014 0.1288 0.1464 0.0349 0.0288 0.0140 0.0134 0.0310 0.0311
No PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) MTLDM PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) MTLDM
@@
Figure 11: Prediction performance and visualization evaluation of the MTLDM model for a precipitation event that occurred on March 15, 2022, at 18:34:00 in the MRMS dataset
CSI Analysis According to Ravuri et al. [3], we evaluate our model’s predictive ability for precipitation intervals of $\geq$ 1 mm/h, and $\geq$ 8.0 mm/h, and we also provide its performance for the 4–8 mm/h interval. Fig. 6 shows the CSI results of sevaral prediction models on the MRMS dataset. The detailed CSI data of various models are listed in Tables II and III. For the precipitation interval of $\geq$ 1 mm/h, the performance of the diffusion model is superior to those of the non-diffusion models, and our proposed MTLDM far outperforms the other methods, including the MTLDM(oa). On average, its CSI is 6% higher than that of the DGMR and 9% higher than that of PySTEPS. For the $\geq$ 4 mm/h region, the CSI of the MTLDM is about 2% higher on average than those of the DGMR and PySTEPS. For the heavy precipitation intervals of 4–8 and $\geq$8 mm/h, the MTLDM also significantly outperforms the other four methods. Fig.7 shows the results of the performance comparison of the six precipitation forecast models for forecasting precipitation events with $\geq$1, 4, 4–8, and $\geq$8 mm/h in the SW datasets according to the CSI, and the detailed data of the models are listed in Table IV and Table V . The MTLDM model performed the best over the entire forecast period, followed by the MTLDM (oa), which had a similar performance. PySTEPS and the DGMR performed similarly, with better results in short-term forecasts, but their performances rapidly degraded over time. At the beginning of the forecast, for the MRMS dataset, the maximum CSI differences between the MTLDM and the other models were 0.16, 0.14, 0.12, and 0.05 for the $\geq$1, $\geq$4, 4–8, and $\geq$8 mm/h scenarios, respectively. For the SW dataset, these differences were approximately 0.11, 0.15, 0.15, and 0.17, respectively. Although the gap has narrowed over time, the MTLDM still maintained the top position.
@@
TABLE VI: Comparison of precipitation CRPS (↓) on the SW dataset. ↓ means the smaller the better
@@
PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa)
1 0.1062 0.0973 0.1036 0.1242 0.0842 0.1073 0.0999 0.1001 0.1063 0.0783 0.0962 0.0924 0.0918 0.0995 0.0694
2 0.1148 0.1108 0.1086 0.1271 0.0922 0.1185 0.1205 0.1062 0.1104 0.0880 0.1099 0.1153 0.0992 0.1042 0.0798
3 0.1221 0.1212 0.1116 0.1302 0.1003 0.1281 0.1385 0.1101 0.1144 0.0965 0.1209 0.1381 0.1039 0.1088 0.0886
4 0.1286 0.1299 0.1155 0.1311 0.1072 0.1364 0.1528 0.1146 0.1160 0.1040 0.1303 0.1575 0.1092 0.1108 0.0969
5 0.1342 0.1428 0.1185 0.1353 0.1123 0.1436 0.1719 0.1181 0.1204 0.1106 0.1383 0.1814 0.1132 0.1156 0.1040
6 0.1391 0.1472 0.1204 0.1384 0.1175 0.1498 0.1816 0.1206 0.1237 0.1160 0.1451 0.1954 0.1162 0.1194 0.1099
7 0.1432 0.1509 0.1228 0.1444 0.1213 0.1550 0.1938 0.1234 0.1285 0.1208 0.1509 0.2199 0.1194 0.1248 0.1152
8 0.1470 0.1468 0.1249 0.1505 0.1245 0.1597 0.1774 0.1258 0.1319 0.1244 0.1561 0.1887 0.1221 0.1285 0.1193
9 0.1505 0.1533 0.1273 0.1532 0.1279 0.1641 0.1924 0.1284 0.1349 0.1282 0.1608 0.2156 0.1250 0.1318 0.1234
10 0.1536 0.1522 0.1287 0.1448 0.1316 0.1680 0.1848 0.1301 0.1314 0.1312 0.1650 0.1989 0.1270 0.1285 0.1267
11 0.1562 0.1572 0.1299 0.1491 0.1335 0.1712 0.2027 0.1316 0.1355 0.1338 0.1685 0.2360 0.1288 0.1329 0.1296
12 0.1588 0.1490 0.1314 0.1540 0.1361 0.1746 0.1843 0.1334 0.1380 0.1364 0.1722 0.2051 0.1309 0.1358 0.1325
13 0.1610 0.1604 0.1327 0.1547 0.1378 0.1775 0.2043 0.1350 0.1391 0.1388 0.1754 0.2359 0.1328 0.1369 0.1352
14 0.1632 0.1506 0.1338 0.1515 0.1399 0.1802 0.1872 0.1363 0.1386 0.1412 0.1783 0.2096 0.1343 0.1366 0.1376
15 0.1648 0.1539 0.1349 0.1498 0.1419 0.1822 0.1880 0.1375 0.1379 0.1429 0.1804 0.2062 0.1354 0.1359 0.1396
16 0.1668 0.1679 0.1375 0.1624 0.1440 0.1846 0.2135 0.1403 0.1477 0.1452 0.1830 0.2452 0.1384 0.1461 0.1420
TABLE VII: Comparison of precipitation CRPS (↓) on the MRMS dataset. ↓ means the smaller the better

### v3

@@ IV-C Diffusion Model Training and Sampling Overview
The diffusion model training involves several steps. Firstly, initial data is sampled from the real data distribution. Then, a time step in the diffusion process is randomly selected, and noise $\epsilon$ is sampled from the standard normal distribution. The difference between the model prediction noise and the actual noise is calculated and minimized. Model parameters are updated via gradient descent, and this process is repeated until the model converges, enhancing the model’s ability to recover the original data from noise. The training is carried out using the PyTorch 1.3.1 framework on 10 A6000 GPU devices under Ubuntu 22.04(Detailed hyperparameter settings are shown in Table I). In the inference phase, the DDPM sampling method is used on a single A800 GPU. It takes approximately 480 seconds to sample a batch of data, which means generating 16 frames of prediction results. For the diffusion model and denoising network, SSLDM and MTLDM have the same structure. The latent space encoder $D_{0}$ and the conditional encoder are also identical. The distinction between the two lies in the need to train different decoders for different decomposed sub-images.
@@
TABLE II: Comparison of Model Sizes
Model Parameters (M)
DGMR 123.17
TRDMv1/v2 48.70
DiffCast 47.07
MTLDM 331.54
MTLDM(oa) 331.54
Figure 6: Comparison of CSI Across Prediction Intervals for Different Precipitation Intensity Thresholds on the SW dataset
Figure 7: Comparison of CRPS Across Prediction Intervals for Different Precipitation Intensity Thresholds on the SW dataset
@@ IV-D Evaluation based on Metrics
No PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) MTLDM PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) MTLDM
@@
1 0.5292 0.5798 0.4686 0.4692 0.6147 0.6243 0.6262 0.2194 0.2621 0.1489 0.1355 0.2579 0.2796 0.2728
2 0.4912 0.5144 0.4448 0.4379 0.5116 0.5670 0.5728 0.1858 0.1949 0.1135 0.1094 0.1718 0.2066 0.2157
3 0.4455 0.4804 0.4249 0.4153 0.4593 0.5272 0.5359 0.1499 0.1585 0.0905 0.0887 0.1358 0.1654 0.1758
4 0.4146 0.4407 0.4083 0.3963 0.4148 0.4943 0.5016 0.1326 0.1292 0.0767 0.0792 0.1122 0.1422 0.1552
5 0.3875 0.4173 0.3931 0.3798 0.3784 0.4677 0.4752 0.1132 0.1090 0.0644 0.0672 0.0925 0.1261 0.1366
6 0.3644 0.3958 0.3805 0.3629 0.3525 0.4439 0.4536 0.1027 0.0947 0.0576 0.0608 0.0809 0.1125 0.1224
7 0.3420 0.3739 0.3648 0.3461 0.3274 0.4229 0.4311 0.0916 0.0827 0.0484 0.0520 0.0718 0.0990 0.1117
8 0.3234 0.3530 0.3524 0.3310 0.3064 0.4015 0.4098 0.0826 0.0727 0.0437 0.0471 0.0664 0.0866 0.0978
9 0.3065 0.3358 0.3408 0.3191 0.2858 0.3831 0.3916 0.0748 0.0642 0.0396 0.0424 0.0595 0.0763 0.0888
10 0.2930 0.3210 0.3306 0.3081 0.2702 0.3690 0.3766 0.0697 0.0573 0.0351 0.0382 0.0546 0.0695 0.080
11 0.2794 0.3074 0.3199 0.2959 0.2545 0.3555 0.3628 0.0643 0.0519 0.0319 0.0345 0.0503 0.0642 0.0753
12 0.2668 0.2945 0.3101 0.2865 0.2402 0.3418 0.3494 0.0584 0.0469 0.0298 0.0327 0.0461 0.0586 0.0682
13 0.2570 0.2830 0.3015 0.2784 0.2297 0.3318 0.3388 0.0542 0.0437 0.0276 0.0306 0.0429 0.0562 0.0650
14 0.2467 0.2719 0.2933 0.2692 0.2159 0.3198 0.3271 0.0522 0.0404 0.0256 0.0273 0.0385 0.0524 0.0601
15 0.2377 0.2617 0.2858 0.2618 0.2053 0.3092 0.3176 0.0494 0.0369 0.0238 0.0260 0.0356 0.0493 0.0564
16 0.2289 0.2512 0.2773 0.2515 0.1960 0.2999 0.3072 0.0462 0.0330 0.0242 0.0260 0.0333 0.0469 0.0537
TABLE IV: Comparison of precipitation CSI (↑) on the MRMS dataset. ↑ means the larger the better
No PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) MTLDM PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) MTLDM
@@
1 0.1880 0.2271 0.1271 0.1163 0.2253 0.2442 0.2355 0.0810 0.1035 0.0411 0.0310 0.0857 0.1048 0.0979
2 0.1565 0.1647 0.0977 0.0950 0.1456 0.1744 0.1818 0.0636 0.0641 0.0318 0.0168 0.0484 0.0660 0.0711
3 0.1257 0.1329 0.0787 0.0771 0.1141 0.1386 0.1471 0.0430 0.0428 0.0234 0.0092 0.0327 0.0457 0.0530
4 0.1103 0.1077 0.0677 0.0700 0.0938 0.1187 0.1273 0.0338 0.0311 0.0184 0.0073 0.0261 0.0366 0.0434
5 0.0946 0.0916 0.0568 0.0597 0.0773 0.1051 0.1138 0.0274 0.0238 0.0144 0.0054 0.0205 0.0286 0.0365
6 0.0854 0.0801 0.0511 0.0539 0.0682 0.0942 0.1030 0.0250 0.0171 0.0133 0.0049 0.0168 0.0243 0.0300
7 0.0760 0.0699 0.0438 0.0472 0.0609 0.0828 0.0919 0.0208 0.0135 0.0107 0.0028 0.0128 0.0189 0.0246
8 0.0690 0.0615 0.0399 0.0429 0.0563 0.0731 0.0819 0.0189 0.0119 0.0076 0.0018 0.0102 0.0152 0.0203
9 0.0628 0.0546 0.0360 0.0384 0.0506 0.0642 0.0747 0.0151 0.0106 0.0073 0.0025 0.0081 0.0122 0.0175
10 0.0586 0.0486 0.0320 0.0350 0.0467 0.0599 0.0681 0.0131 0.0080 0.0063 0.0019 0.0065 0.0091 0.0143
11 0.0540 0.0444 0.0290 0.0313 0.0434 0.0549 0.0641 0.0122 0.0067 0.0060 0.0014 0.0052 0.0070 0.0117
12 0.0497 0.0404 0.0276 0.0302 0.0400 0.0506 0.0591 0.0109 0.0064 0.0048 0.0015 0.0037 0.0056 0.0105
13 0.0462 0.0379 0.0259 0.0285 0.0373 0.0493 0.0560 0.0095 0.0052 0.0038 0.0011 0.0037 0.0059 0.0086
14 0.0448 0.0348 0.0238 0.0257 0.0337 0.0462 0.0520 0.0093 0.0050 0.0038 0.0008 0.0032 0.0048 0.0079
15 0.0422 0.0318 0.0221 0.0241 0.0314 0.0432 0.0487 0.0079 0.0051 0.0039 0.0009 0.0029 0.0055 0.0076
16 0.0393 0.0282 0.0222 0.0239 0.0295 0.0407 0.0467 0.0071 0.0048 0.0041 0.0013 0.0022 0.0048 0.0078
No PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) MTLDM PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) MTLDM
@@
1 0.4235 0.3773 0.2889 0.2655 0.3735 0.3934 0.4663 0.1882 0.1657 0.0928 0.0740 0.1475 0.1602 0.2642
2 0.3765 0.2878 0.2550 0.2443 0.3084 0.3497 0.4148 0.1557 0.1132 0.0729 0.0634 0.1133 0.1331 0.2067
3 0.3386 0.2724 0.2322 0.2267 0.2645 0.3155 0.3737 0.1320 0.1105 0.0602 0.0559 0.0904 0.1134 0.1649
4 0.3084 0.2368 0.2137 0.2117 0.2307 0.2855 0.3318 0.1135 0.0871 0.0512 0.0507 0.0724 0.0965 0.1337
5 0.2821 0.2090 0.1962 0.1971 0.2014 0.2588 0.3029 0.0993 0.0647 0.0438 0.0423 0.0627 0.0829 0.1114
6 0.2596 0.1859 0.1821 0.1840 0.1784 0.2365 0.2761 0.0865 0.0527 0.0375 0.0379 0.0522 0.0717 0.0954
7 0.2420 0.1791 0.1701 0.1711 0.1607 0.2188 0.2564 0.0787 0.0528 0.0330 0.0331 0.0454 0.0655 0.0836
8 0.2265 0.1634 0.1602 0.1614 0.1454 0.2041 0.2371 0.0710 0.0441 0.0291 0.0296 0.0395 0.0576 0.0728
9 0.2116 0.1586 0.1508 0.1518 0.1310 0.1898 0.2206 0.0639 0.0393 0.0261 0.0259 0.0347 0.0518 0.0639
10 0.1991 0.1507 0.1414 0.1437 0.1183 0.1781 0.2055 0.0593 0.0379 0.0240 0.0226 0.0305 0.0481 0.0561
11 0.1879 0.1408 0.1342 0.1356 0.1062 0.1672 0.1934 0.0544 0.0362 0.0210 0.0218 0.0268 0.0433 0.0501
12 0.1778 0.1279 0.1260 0.1286 0.0967 0.1573 0.1816 0.0499 0.0359 0.0183 0.0194 0.0238 0.0403 0.0446
13 0.1685 0.1371 0.1191 0.1210 0.0876 0.1489 0.1712 0.0458 0.0382 0.0164 0.0170 0.0214 0.0383 0.0404
14 0.1594 0.1142 0.1117 0.1136 0.0786 0.1418 0.1619 0.0409 0.0320 0.0151 0.0152 0.0197 0.0346 0.0367
15 0.1519 0.1164 0.1059 0.1069 0.0710 0.1350 0.1546 0.0376 0.0258 0.0149 0.0141 0.0172 0.0323 0.0339
16 0.1444 0.1227 0.1001 0.1014 0.0634 0.1288 0.1464 0.0349 0.0288 0.0140 0.0134 0.0148 0.0310 0.0311
TABLE VI: Comparison of precipitation CSI (↑) on the SW dataset. ↑ means the larger the better
No PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) MTLDM PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) MTLDM
@@
Figure 10: Prediction performance and visualization evaluation of the MTLDM model for a precipitation event that occurred on April 10, 2022, at 15:16:00 in the MRMS dataset
CSI Analysis According to Ravuri et al. [3], we evaluate our model’s predictive ability for precipitation intervals of $\geq$ 1 mm/h, and $\geq$ 8.0 mm/h, and we also provide its performance for the 4–8 mm/h interval. Fig. 5 shows the CSI results of sevaral prediction models on the MRMS dataset. The detailed CSI indicator data of the prediction results of each model on the MRMS dataset are listed in TableIII and TableIV. For the precipitation interval of $\geq$ 1 mm/h, the performance of the diffusion model is superior to those of the non-diffusion models, and our proposed MTLDM far outperforms the other methods, including the MTLDM(oa). On average, its CSI is 13% higher than DGMR, 23% higher than PySTEPS, and 26.5% higher than Diffcast. For the $\geq$ 4 mm/h region, the CSI of the MTLDM is about 20% higher on average than those of the DGMR and PySTEPS and also outperforms DiffCast. For the heavy precipitation intervals of 4-8 and $\geq$8 mm/h, the MTLDM also significantly outperforms the other four methods. Fig.6 shows the results of the performance comparison of the six precipitation forecast models for forecasting precipitation events with $\geq$1, 4, 4-8, and $\geq$8 mm/h in the SW datasets according to the CSI, and the detailed data of the models are listed in Table V and Table VI . The MTLDM model has the best overall performance over the entire forecast period, followed by MTLDM (oa) with similar performance. DiffCast performed better than TRDM models but not as well as MTLDM. PySTEPS and the DGMR performed similarly, with better results in short-term forecasts, but their performances rapidly degraded over time. At the beginning of the forecast, for the MRMS dataset, the maximum CSI differences between the MTLDM and the other models were 0.16, 0.14, 0.12, and 0.05 for the $\geq$1, $\geq$4, 4–8, and $\geq$8 mm/h scenarios, respectively. For the SW dataset, these differences were approximately 0.11, 0.15, 0.15, and 0.17, respectively. Although the gap has narrowed over time, the MTLDM still maintained the top position.
@@
TABLE VII: Comparison of precipitation CRPS (↓) on the SW dataset. ↓ means the smaller the better
@@
PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa)
1 0.1062 0.0973 0.1036 0.1242 0.0887 0.0842 0.1073 0.0999 0.1001 0.1063 0.0843 0.0783 0.0962 0.0924 0.0918 0.0995 0.0776 0.0694
2 0.1148 0.1108 0.1086 0.1271 0.1092 0.0922 0.1185 0.1205 0.1062 0.1104 0.1064 0.0880 0.1099 0.1153 0.0992 0.1042 0.1011 0.0798
3 0.1221 0.1212 0.1116 0.1302 0.1251 0.1003 0.1281 0.1385 0.1101 0.1144 0.1233 0.0965 0.1209 0.1381 0.1039 0.1088 0.1194 0.0886
4 0.1286 0.1299 0.1155 0.1311 0.1392 0.1072 0.1364 0.1528 0.1146 0.1160 0.1381 0.1040 0.1303 0.1575 0.1092 0.1108 0.1353 0.0969
5 0.1342 0.1428 0.1185 0.1353 0.1497 0.1123 0.1436 0.1719 0.1181 0.1204 0.1494 0.1106 0.1383 0.1814 0.1132 0.1156 0.1478 0.1040
6 0.1391 0.1472 0.1204 0.1384 0.1522 0.1175 0.1498 0.1816 0.1206 0.1237 0.1522 0.1160 0.1451 0.1954 0.1162 0.1194 0.1509 0.1099
7 0.1432 0.1509 0.1228 0.1444 0.1569 0.1213 0.1550 0.1938 0.1234 0.1285 0.1571 0.1208 0.1509 0.2199 0.1194 0.1248 0.1564 0.1152
8 0.1470 0.1468 0.1249 0.1505 0.1532 0.1245 0.1597 0.1774 0.1258 0.1319 0.1612 0.1244 0.1561 0.1887 0.1221 0.1285 0.1609 0.1193
9 0.1505 0.1533 0.1273 0.1634 0.1607 0.1279 0.1641 0.1924 0.1284 0.1349 0.1641 0.1282 0.1608 0.2156 0.1250 0.1318 0.1640 0.1234
10 0.1536 0.1522 0.1287 0.1448 0.1655 0.1316 0.1680 0.1848 0.1301 0.1314 0.1665 0.1312 0.1650 0.1989 0.1270 0.1285 0.1669 0.1267
11 0.1562 0.1572 0.1299 0.1491 0.1630 0.1335 0.1712 0.2027 0.1316 0.1355 0.1640 0.1338 0.1685 0.2360 0.1288 0.1329 0.1645 0.1296
12 0.1588 0.1490 0.1314 0.1540 0.1636 0.1361 0.1746 0.1843 0.1334 0.1380 0.1649 0.1364 0.1722 0.2051 0.1309 0.1358 0.1656 0.1325
13 0.1610 0.1604 0.1327 0.1547 0.1635 0.1378 0.1775 0.2043 0.1350 0.1391 0.1649 0.1388 0.1754 0.2359 0.1328 0.1369 0.1657 0.1352
14 0.1632 0.1506 0.1338 0.1515 0.1634 0.1399 0.1802 0.1872 0.1363 0.1386 0.1649 0.1412 0.1783 0.2096 0.1343 0.1366 0.1659 0.1376
15 0.1648 0.1539 0.1349 0.1498 0.1632 0.1419 0.1822 0.1880 0.1375 0.1379 0.1648 0.1429 0.1804 0.2062 0.1354 0.1359 0.1660 0.1396
16 0.1668 0.1679 0.1375 0.1624 0.1600 0.1440 0.1846 0.2135 0.1403 0.1477 0.1661 0.1452 0.1830 0.2452 0.1384 0.1461 0.1629 0.1420
TABLE VIII: Comparison of precipitation CRPS (↓) on the MRMS dataset. ↓ means the smaller the better